Water falling from a 50 m high fall is to be used for generating electric energy. If 1.8 × 105 kg of water falls per hour and half the gravitational potential energy can be converted into electric energy, how many 100 W lamps can be lit?

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. . h = 50m, m = 1.8 × 10^5 kg/hr, P = 100 watt, P.E. = mgh = 1.8 × 10^5 × 9.8 × 50 = 882 × 10^5 J/hr Because, half the potential energy is converted into electricity, Electrical energy ½ P.E. = 441 × 10^5 J/hr So, power in watt (J/sec) is given by = 441 * 10^5/3600 ∴ number of 100 W lamps, that can be lit 441 * 10^5/3600 * 100 = 122.5 = 122